# Toward Autonomous Neural VMC: An Energy-Variance Convergence Criterion for Quantum Systems

### v1

@@ 3.1.2 Results and analysis of multiple physical systems
$\hat{H}_{\mathrm{HO}}=-\frac{1}{2}\nabla^{2}+\frac{1}{2}r^{2},$ (13)
$\hat{H}_{\mathrm{Hyd}}=-\frac{1}{2}\nabla^{2}-\frac{1}{r}.$ (14)
@@
The relationship between the final energy variance and relative error across all test cases is presented in Fig. 1. While a moderate correlation is observed on a logarithmic scale (Pearson correlation coefficient = 0.598), the most significant finding emerges from the empirical distribution of the data points: all systems achieving an energy variance below $1\times 10^{-3}$ exhibit relative errors under 1%. This consistency across systems—from simple harmonic potentials to complex hadronic interactions—demonstrates the broad applicability of our variance-based criterion. We therefore propose $1\times 10^{-3}$ as a universal threshold of energy variance for indicating preliminary convergence across diverse quantum systems.This criterion is inherently scalable: demanding higher precision simply requires lowering the variance threshold accordingly.
@@
The distinct advantage of the energy-variance criterion is further highlighted in the case of the $\chi_{c}$ particle. As illustrated in Fig. 2c, the energy curve reaches an apparent plateau, suggesting convergence by a traditional standard. However, the energy variance continues its steady decline, guiding the optimization to a solution of significantly higher accuracy. This demonstrates the criterion’s unique ability to identify and overcome deceptive pseudo-convergence, ensuring that the final result represents the best possible solution within the representational capacity of the neural network.
@@ 3.2 Parameter scanning of multiple systems
The initial stage of physical research often prioritizes understanding how system properties evolve with external conditions or internal parameters, over pursuing the ultimate precision for a single configuration. A reliable, hands-off convergence criterion is the cornerstone of such large-scale, automated computational campaigns. Leveraging the energy-variance criterion, we have developed a lightweight and general-purpose solver whose rapid convergence and low computational footprint make it ideally suited for batch parameter scanning. This capability allows us to efficiently map the “phase diagrams” of ground-state properties, providing an intuitive understanding of parametric dependencies and crucial guidance for subsequent focused studies.
@@ 3.2.2 Compression of the hydrogen atom in the strong magnetic field
Plot the curve of the system energy versus magnetic induction intensity and fit it with a quadratic polynomial as shown in Fig. 5. The ground state energy increases monotonically with the magnetic field strength, rising from -0.4956 a.u. at $B=0.1$ a.u. to -0.327 a.u. at $B=1.0$ a.u. This trend is in full agreement with theoretical expectations and is primarily attributed to the positive energy contribution from the diamagnetic term in the Hamiltonian.
@@
The changes in the wave function are relatively subtle shown as Fig. 6. By employing a logarithmic scale for visualization, it becomes apparent that the wave function is compressed in the xy-plane (perpendicular to the magnetic field) as the field strength increases. Furthermore, we observe a slight, systematic contraction along the z-axis. This indicates that under a strong magnetic field, the electron cloud of the hydrogen atom undergoes a comprehensive three-dimensional redistribution, rather than being affected only in the plane perpendicular to the field.
@@ 3.2.3 Three-body quantum dot in a anisotropic harmonic oscillator trap
$g_{ij}(r)=\frac{1}{N}\left\langle\delta(r-|\mathbf{r}_{i}-\mathbf{r}_{j}|)\right\rangle,$ (21)
where N is normalization factor. It describes the probability density of this particle pair occurring at different inter-particle distances.
@@
3.3 Discussion on the robustness and limitations of the energy-variance convergence criterion
While the energy-variance criterion is theoretically rigorous, its practical implementation requires careful consideration of numerical and physical complexities. A key concern for any convergence criterion is the occurrence of false positives—where the criterion is met, but the solution is physically incorrect. In our experiments, a salient example was observed in the weakly-coupled double-well potential at Sec. 3.2.1.
@@ 3.3 Discussion on the robustness and limitations of the energy-variance convergence criterion
Our current framework employs several strategies to enhance robustness, including large walker counts and monitoring the moving average of the variance. For future work, tackling near-degeneracy may require incorporating symmetry projections or employing optimization techniques specifically designed to escape shallow local minima. Nevertheless, across the vast majority of our tests, the energy-variance criterion proved to be a remarkably reliable and universal indicator of convergence, successfully automating the parameter scans presented herein.
@@ 4 Conclusion
In this work, we have systematically investigated and validated the energy variance as a robust and practical convergence criterion for neural wave function optimization in variational Monte Carlo. Our central contribution is the translation of a core quantum-mechanical principle—that an eigenstate exhibits zero variance—into an automated and computationally tractable tool.
We demonstrated the reliability and universality of our criterion through testing systems spanning different interaction types (harmonic, Coulomb, hadronic). The empirical finding that an energy variance below $1\times 10^{-3}$ guarantees a relative error under 1% provides a clear and accessible standard for the community. This objective threshold served as the foundation for automating parameter scans. By integrating the criterion into a lightweight neural solver, we enabled fully autonomous workflows, significantly accelerating the preliminary physical verification of quantum Hamiltonians.
The computational efficiency of our framework, achieving accurate results with modest resources, underscores its advantages in accessibility and automation. By offering a quantitative, hands-off convergence signal, our method lowers the barrier to reliable quantum-mechanical calculations, establishing it as an ideal platform for high-throughput preliminary studies.
Looking ahead, this work opens several promising avenues. The energy-variance criterion can be directly extended to more complex systems, including fermionic matter and frustrated lattice models. Methodologically, the solver’s scalability allows for architectural specialization, such as incorporating equivariant networks to build in physical symmetries for enhanced efficiency. A particularly compelling challenge lies in extending this framework to time-dependent Hamiltonians, potentially enabling the study of quantum dynamics through real-time evolution.
@@ Acknowledgements
Appendix A
This appendix provides visualizations of the wave functions for the benchmark quantum systems discussed in the Sec. 3.1.2, serving as further validation of the physical correctness of our computational results. All wave functions were optimized by our solver under the guidance of the energy-variance convergence criterion. Their morphologies are in precise agreement with the theoretical expectations for their respective systems.
@@
The reduced radial probability density shown as Fig. 14, $u(r)=rR(r)$, for the four charmonium states ($\eta_{c}$, $J/\psi$, $h_{c}$, $\chi_{c}$). Their shapes and spatial extents correctly reflect the physical characteristics under different orbital angular momenta and spin configurations, confirming our method’s capability for handling complex many-body interaction potentials.

### v2

@@ 3.1.2 Results and analysis of multiple physical systems
$\hat{H}_{\mathrm{HO}}=-\frac{1}{2m}\nabla^{2}+\frac{1}{2}m\omega^{2}r^{2},$ (12)
$\hat{H}_{\mathrm{Hyd}}=-\frac{1}{2}\nabla^{2}-\frac{1}{r},$ (13)
where $m=\omega^{2}=1.0$ in the harmonic oscillator of our setting.
@@
The relationship between the final energy variance and relative error across all test cases is presented in Fig. 1. While a moderate correlation is observed on a logarithmic scale (Pearson correlation coefficient = 0.598), the most significant finding emerges from the empirical distribution of the data points: all systems achieving an energy variance below $1\times 10^{-3}$ exhibit relative errors under 1%. This consistency across systems—from simple harmonic potentials to complex hadronic interactions—demonstrates the broad applicability of our variance-based criterion. We therefore use $1\times 10^{-3}$ as a threshold of energy variance for indicating preliminary convergence across diverse quantum systems.This criterion is inherently scalable: demanding higher precision simply requires lowering the variance threshold accordingly.
@@
The distinct advantage of the energy variance criterion is further highlighted in the case of the $\chi_{c}$ particle. As illustrated in Fig. 2c, the energy curve reaches an apparent plateau, suggesting convergence by a traditional standard. However, the energy variance continues its steady decline, guiding the optimization to a solution of significantly higher accuracy. This demonstrates the criterion’s unique ability to identify and overcome deceptive pseudo-convergence, ensuring that the final result represents the best possible solution within the representational capacity of the neural network.
@@ 3.2 Parameter scanning of multiple systems
The initial stage of physical research often prioritizes understanding how system properties evolve with external conditions or internal parameters, over pursuing the ultimate precision for a single configuration. A reliable, hands-off convergence criterion is the cornerstone of such large-scale, automated computational campaigns. Leveraging the energy variance criterion, we have developed a lightweight and general-purpose solver whose rapid convergence and low computational footprint make it ideally suited for batch parameter scanning. This capability allows us to efficiently map the “phase diagrams” of ground-state properties, providing an intuitive understanding of parametric dependencies and crucial guidance for subsequent focused studies.
@@ 3.2.2 Compression of the hydrogen atom in the strong magnetic field
Plot the curve of the system energy versus magnetic induction intensity and fit it with a quadratic polynomial as shown in Fig. 5. The ground state energy increases monotonically with the magnetic field strength, rising from -0.4956 a.u. at $B=0.1$ a.u. to -0.327 a.u. at $B=1.0$ a.u.. This trend aligns well with theoretical predictions and is primarily attributed to the positive energy contribution from the diamagnetic term in the Hamiltonian.
@@
The changes in the wave function are relatively subtle shown as Fig. 6. By employing a logarithmic scale for visualization, it becomes apparent that the electron density exhibits significant compression in the xy-plane (perpendicular to the magnetic field) as the field strength increases. Furthermore, we observe a slight, systematic contraction along the z-axis. This indicates that under a strong magnetic field, the electron cloud of the hydrogen atom undergoes a comprehensive three-dimensional redistribution, rather than being affected only in the plane perpendicular to the field.
@@ 3.2.3 Three-body quantum dot in a anisotropic harmonic oscillator trap
$g_{ij}(r)=C\left\langle\delta(r-|\mathbf{r}_{i}-\mathbf{r}_{j}|)\right\rangle,$ (20)
where $C$ is normalization factor. It describes the probability density of this particle pair occurring at different inter-particle distances.
@@
3.3 Discussion on the robustness and limitations of the energy variance convergence criterion
While the energy variance criterion is theoretically rigorous, its practical implementation requires careful consideration of numerical and physical complexities. A key concern for any convergence criterion is the occurrence of false positives—where the criterion is met, but the solution is physically incorrect. In our experiments, a salient example was observed in the weakly-coupled double-well potential at Sec. 3.2.1.
@@ 3.3 Discussion on the robustness and limitations of the energy variance convergence criterion
This instance highlights a fundamental trade-off between different convergence philosophies. While powerful frameworks like NetKet[27] employ sophisticated statistical diagnostics like the Gelman-Rubin statistic to ensure robust sampling—enabling them to tackle extremely challenging systems—our approach prioritizes physical interpretability and computational agility. The energy variance provides a direct, quantum-mechanical measure of wave function quality, and its implementation requires only a single optimization run. This makes our paradigm particularly suited for the rapid preliminary investigation of parameterized Hamiltonians. The two approaches are complementary: our variance criterion could, in future work, be integrated as a high-level physical validation step within more complex frameworks that use statistical diagnostics for low-level sampling control.
@@ 4 Conclusion
In this work, we have established the energy variance as a robust and practical convergence criterion for neural wave function optimization in variational Monte Carlo. By translating the fundamental quantum-mechanical principle—that eigenstates exhibit zero variance—into an automated computational framework, we have developed a lightweight neural solver that enables efficient and reliable calculations for quantum systems. A theoretical analysis of the variance landscape further clarifies the domain of applicability of our criterion, while comprehensive numerical validations confirm its effectiveness across diverse systems. Our approach consistently delivers accurate results with modest computational resources, underscoring its strengths in accessibility and scalability.
Across a range of non-nodal ground-state systems—including harmonic oscillators, hydrogen atoms, and charmonium—variance minimization proves both effective and reliable. We further propose an empirical variance threshold of $10^{-3}$ as a convergence standard, which facilitates automated parameter scans and significantly accelerates the preliminary physical verification process. It should be noted that the precise value of this threshold may vary with system size and complexity. The successful application of the variance as a convergence diagnostic in excited states further demonstrates that, although variance minimization may not always serve as a stable optimization parameter in nodal structures, the variance itself remains a valuable indicator of convergence.
Looking forward, future work will focus on extending the applicability of our framework by developing technical strategies to mitigate singularities, thereby effectively generalizing the energy-variance criterion to nodal systems. This advancement will further broaden the utility of our approach to strongly correlated and complex fermionic systems.
@@ Acknowledgements
Data Availability Statement
The data that support the findings of this study are available within the article and its Supplementary Material. The source code used in this study will be made publicly available on GitHub upon publication.
Appendix A Wave function visualizations
This appendix provides visualizations of the wave functions for the benchmark quantum systems discussed in the Sec. 3.1.2, serving as further validation of the physical correctness of our computational results. All wave functions were optimized by our solver under the guidance of the energy-variance convergence criterion. Their morphologies are in agreement with the theoretical expectations for their respective systems.
@@ Appendix A Wave function visualizations
The reduced radial probability density shown as Fig. 14, $u(r)=rR(r)$, for the four charmonium states ($\eta_{c}$, $J/\psi$, $h_{c}$, $\chi_{c}$). Their shapes and spatial extents correctly reflect the physical characteristics under different orbital angular momenta and spin configurations, confirming our method’s capability for handling complex interaction potentials.